OpenStudy (anonymous):

PLEASE HELP Let C(q) represent the total cost of producing q items. Then C′(q) gives the marginal cost in dollars per item. Suppose a company determines that C(13)=2800 and C′(13)=73. Estimate the total cost of producing the following number of items. Your answers should be whole numbers. a) 12 items

OpenStudy (anonymous):

You could use the tangent line approximation: \[ \frac{L(q)-C(q_0)}{q-q_0}=C'(q_0) \implies L(q)=(q-q_0)C'(q_0)+C(q_0) \]Basically, \(L(q)\) is a linear approximation of \(C(q)\) \(q_0\) is the point for which we have the derivative (in this case \(13\)) \(q\) is the point we are trying to approximate at (in this case \(12\))

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so how do we plug into that

OpenStudy (anonymous):

Use the right most equality I gave you, and plug in the values for \(q\) and \(q_0\) which I specified.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and

OpenStudy (anonymous):

@wio what'd u get